# Your Title

### v1

@@ Abstract
3D Gaussian Splatting (3DGS) has emerged as a high-fidelity and efficient paradigm for online free-viewpoint video (FVV) reconstruction, offering viewers rapid responsiveness and immersive experiences. However, existing online methods face challenge in prohibitive storage requirements primarily due to point-wise modeling that fails to exploit the motion properties. To address this limitation, we propose a novel Compact Gaussian Streaming (ComGS) framework, leveraging the locality and consistency of motion in dynamic scene, that models object-consistent Gaussian point motion through keypoint-driven motion representation. By transmitting only the keypoint attributes, this framework provides a more storage-efficient solution. Specifically, we first identify a sparse set of motion-sensitive keypoints localized within motion regions using a viewspace gradient difference strategy. Equipped with these keypoints, we propose an adaptive motion-driven mechanism that predicts a spatial influence field for propagating keypoint motion to neighboring Gaussian points with similar motion. Moreover, ComGS adopts an error-aware correction strategy for key frame reconstruction that selectively refines erroneous regions and mitigates error accumulation without unnecessary overhead. Overall, ComGS achieves a remarkable storage reduction of over 159 $\times$ compared to 3DGStream and 14 $\times$ compared to the SOTA method QUEEN, while maintaining competitive visual fidelity and rendering speed. Our code will be released.
Figure 1: Left: Experimental results on N3DV dataset [14] showcase the effectiveness of our method, which reduces the storage requirement of 3DGStream [13] by 159 $\times$, with enhanced visual quality. Right: Comparison with existing methods in storage and reconstruction fidelity. Hollow circles denote offline methods, while solid circles represent online methods.
@@ 1 Introduction
Reconstructing free-viewpoint video (FVV) from multi-view videos captured by cameras with known poses has attracted growing interest in the field of computer vision and graphics. FVV exhibits great potential as a next-generation visual medium that enables immersive and interactive experiences, with broad application in virtual and augmented reality (VR/AR) applications [26].
@@
Recently, 3D Gaussian Splatting (3DGS) has become a promising method for FVV reconstruction, due to its significant advancements in real-time rendering and high-fidelity view synthesis. These approaches typically fall into two categories: 1) incorporating temporal function into Gaussian primitives and optimizing directly [19, 34, 12], and 2) applying a deformation field to capture the spatio-temporal transformations of canonical Gaussians [18, 11, 37, 38, 39]. While these FVV reconstructions accurately represent dynamic scenes, they are trained in an offline manner and require transmitting the full set of reconstructed parameters prior to rendering.
@@
In contrast, by enabling per-frame training and progressive transmission, online FVV reconstruction allows immediate playback without the overhead of full-scene preloading. As a pioneer work, 3DGStream [13] extends 3DGS to online FVV reconstruction using InstantNGP [4] to model the geometric transformation frame-by-frame. While achieving impressive rendering speed, its structural constraint hinders the volumetric representation performance and degrades the visual quality. Building on this paradigm, subsequent works [32, 33] enhance model expressiveness through explicitly optimizing Gaussian attribute residuals, achieving competitive synthesis quality and higher robustness. However, the storage demands of these methods remain prohibitively high for real-time transmission, with reconstructed data typically exceeding 20 MB per second.
@@
In this paper, we aim to design a storage-efficient solution for FVV streaming that minimizes bandwidth requirements and enables real-time transmission. In online FVV reconstruction, since dynamic scenes contain a large proportion of static regions, the key to efficient reconstruction lies in motion modeling. Our first insight, therefore, is to only model the Gaussian attribute residuals in the motion regions, which eliminates the unnecessary updates in static regions. Building on motion modeling, we note that scene motion tends to be consistent, where Gaussian points associated with the same object typically exhibit the same or similar motion in dynamic scene representation. Our second insight, based on this observation, is to use a shared motion representation to model the attribute residuals with similar motion. This contrasts with existing online methods [13, 32] that utilize point-wise strategy to update the attribute residuals in motion regions, and the result is motion redundancy elimination and more compact storage. Lastly, we exploit a key frame fine-tune strategy to handle the error accumulation brought by non-rigid motion and novel objects emergence.
@@
Specifically, to accomplish this, we propose a Compact Gaussian Streaming (ComGS) framework that leverages a set of keypoints ($=200$), significantly fewer than the full set of Gaussian points ($\approx 200K$), to holistically model motion regions at each timestep. ComGS begins with a motion-sensitive keypoints selection through a viewspace gradient difference strategy. This ensures that the selected keypoints are accurately positioned within motion regions and prevents redundant or incorrect modeling of static areas. Subsequently, we design an adaptive motion-driven mechanism that defines a keypoint-specific spatial influence field, with which neighboring Gaussian points can share the motion of the keypoint. Unlike conventional k-nearest neighbor (KNN) methods [43, 44], the spatial influence field can accommodate the complexity and variability of motion structure in dynamic scenes, so that keypoints can more accurately drive the motion of the surrounding region. Finally, to mitigate error accumulation in a compact and effective manner, we propose an error-aware correction strategy for key frame reconstruction that selectively updates only those Gaussians with reconstruction errors.
@@
  • Experiments on two benchmark datasets show that the effectiveness of our method and its individual components. Our method achieves a compression ratio of 159×159$\times$159 × over the 3DGStream and 14×14$\times$14 × over state-of-the-art model QUEEN, enabling real-time transmission while preserving competitive reconstruction quality and rendering speed.
@@ 2.1 Dynamic Gaussian Splatting
Recently, 3D Gaussian Splatting (3DGS) [9, 10, 20, 21, 22, 23] has attracted great attention in Free-viewpoint Video (FVV) reconstruction for its high photorealistic performance and real-time rendering speed. Several works [19, 34, 12] expand temporal variation as a function and optimize directly for modeling Gaussian attributes across frames. For instance, 4D Gaussian Splatting [19] incorporates time-conditioned 3D Gaussians and auxiliary components into 4D Gaussians, while ST-GS [12] models the transformation of structural attributes and opacity as a temporal function to represent scene motions. These time variant-based methods achieve superior rendering efficiency, but suffer from prohibitive storage requirements. Other works [11, 39, 40, 41, 42] employ vanilla 3D Gaussians as a canonical space and a deformation field to represent the dynamic scene. In this category, 4D-GS [11] utilizes hexplanes [15], six orthogonal planes, as latent embeddings and deliver them into a small MLP to deform temporal transformation of Gaussian points, achieving efficient computational complexity and lightweight storage requirement. Building upon this, GD-GS [39] further improves scene modeling accuracy by incorporating geometric priors, which provides a more structured and precise representation of dynamic scene. Among them, both SC-GS [43] and SP-GS [44] adopt sparse control points to control scene motion using a k-nearest neighbor (KNN) [45] strategy for motion modeling. While these methods achieve notable improvements in computational efficiency and rendering speed, they are designed for offline FVV reconstruction and do not support frame-by-frame delivering. Additionally, motion-insensitive control point selection and scale-agnostic KNN motion modeling lead to redundant representation of static regions and reduced deformation accuracy in dynamic scenes. Our online method addresses these limitations by selecting keypoints from motion regions at each timestep and modeling motion with awareness of local motion scales, which enables more accurate and efficient modeling of online FVV.
@@ 2.2 Online Free-Viewpoint Video Reconstruction
Compared to the offline methods, online reconstruction enables FVV to be incrementally trained and transmitted in a per-frame manner, which allows users to preview or interact immediately with the video content. Leveraging the high-fidelity view synthesis capabilities of Neural Radiance Field (NeRF) [1, 2, 3, 4, 5, 6, 7, 8], a set of studies have explored NeRF-based methods [25, 46, 47, 36, 29] for online FVV reconstruction, such as StreamRF [25], VideoRF [46] and TeTriRF [36]. Despite advanced visual quality, NeRF-based methods are hindered by their limited rendering speeding of implicit structure, which limits their practical applications.
@@
With the utilization of 3DGS [9], 3DGStream [13] introduces a hash-based MLP to encode the position and rotation transformation of Gaussian points at each frame, and designs an adaptive Gaussians addition strategy for novel objects across frames. Based on this paradigm, QUEEN [32] proposes a Gaussian residual-based framework for model expressiveness enhancement and a learned quantization-sparsity framework for residuals compression. HiCoM [33] designs a hierarchical coherent motion mechanism to effectively capture and represent scene motion for fast and accurate training. To deploy into mobile device, V3 [28] presents a novel approach that compresses Gaussian attributes as a 2D video to facilitate hardware video codecs. IGS [48] proposes a generalized anchor-driven Gaussian motion network that learns residuals with a singe step, achieving a significant improvement of training speed. Nevertheless, these methods face challenge in real-time transmission, due to their substantial storage requirements. This overhead mainly stems from redundant updates of both static Gaussian points across frames, as well as repeated modeling of Gaussian points with similar motion. Our study exploits the locality and consistency of motion by leveraging motion-sensitive keypoints to adaptively drive motion regions, and this avoids redundant storage and transmission.
@@
Figure 2: The overall pipeline of ComGS framework. (a) The reconstruction process starts from the first frame initialized using vanilla 3DGS [9]. Subsequent frames are organized into groups of frames (GoFs). For non-key frames, (b) we begins with a motion-sensitive keypoint selection using a viewspace gradient difference strategy, (c) and utilizes an adaptive motion-driven mechanism to control neighboring points motion. For key frames, (d) an error-aware correction strategy is introduced to mitigate the error accumulation across frames.
@@ 3.1 Preliminary
3DGS [9] models a 3D scene as a large amount of anisotropic 3D Gaussian points in world space as an explicit representation. The central position and geometric shape of each Gaussian point in world space are defined by a mean vector $\mu$ and covariance matrix $\Sigma$, mathematically represented as:
$G(x)=\text{exp}({-\frac{1}{2}(\textbf{x}-\mu)^{T}\Sigma^{-1}(\textbf{x}-\mu))}$ (1)
For differentiable optimization, the covariance matrix $\Sigma$ is decoupled into a scaling matrix $S$ and a rotation matrix $R$. Each Gaussian point is characterized by its color $c_{i}$ and opacity $\alpha$.
During rendering, the Gaussian points are initially projected into viewing plane, and the color of each pixel $C$ can be obtained by $\alpha$-blending:
$C=\sum_{i\in n}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$ (2)
where $n$ is the number of Gaussian points contributing to this pixel.
@@ 3.2 Motion-Sensitive Keypoint Selection
Thus, inspired by [32], we propose a motion-sensitive keypoint selection based on viewspace gradient difference (Fig. 2 (b)). The core idea is to identify the dynamic Gaussian by the gradient change of rendering loss in inter-frames, and based on the gradient values, the $k$ Gaussian points with the largest gradient are selected as keypoints. Specifically, following the gradient computation in 3DGS [9], we compute gradients using the previous Gaussian positions $p_{t-1}$, the rendered images $R_{t-1}$, the reconstruction loss $\mathcal{L}{{}_{recon}}$, and the ground-truth images $GT_{t-1}$ and $GT_{t}$:
$\mathcal{G}_{t-1}=\frac{\partial\mathcal{L}_{recon}^{t-1}}{\partial p_{t-1}},\ \mathcal{L}_{recon}^{t-1}=\mathcal{L}_{recon}(R_{t-1},GT_{t-1})$ (3)
$\mathcal{G}_{t}=\frac{\partial\mathcal{L}_{recon}^{t}}{\partial p_{t-1}},\ \mathcal{L}_{recon}^{t}=\mathcal{L}_{recon}(R_{t-1},GT_{t})$ (4)
Dynamic significance scores $\Delta\mathcal{G}_{t}\in\mathbb{R}^{N}$ ($N$ is the number of Gaussians) were calculated by means of absolute values of gradient differences:

### v2

# Motion Matters: Compact Gaussian Streaming for Free-Viewpoint Video Reconstruction

@@ Abstract
3D Gaussian Splatting (3DGS) has emerged as a high-fidelity and efficient paradigm for online free-viewpoint video (FVV) reconstruction, offering viewers rapid responsiveness and immersive experiences. However, existing online methods face challenge in prohibitive storage requirements primarily due to point-wise modeling that fails to exploit the motion properties. To address this limitation, we propose a novel Compact Gaussian Streaming (ComGS) framework, leveraging the locality and consistency of motion in dynamic scene, that models object-consistent Gaussian point motion through keypoint-driven motion representation. By transmitting only the keypoint attributes, this framework provides a more storage-efficient solution. Specifically, we first identify a sparse set of motion-sensitive keypoints localized within motion regions using a viewspace gradient difference strategy. Equipped with these keypoints, we propose an adaptive motion-driven mechanism that predicts a spatial influence field for propagating keypoint motion to neighboring Gaussian points with similar motion. Moreover, ComGS adopts an error-aware correction strategy for key frame reconstruction that selectively refines erroneous regions and mitigates error accumulation without unnecessary overhead. Overall, ComGS achieves a remarkable storage reduction of over 159 $\times$ compared to 3DGStream and 14 $\times$ compared to the SOTA method QUEEN, while maintaining competitive visual fidelity and rendering speed. Project page: https://chenjiacong-1005.github.io/ComGS/.
Figure 1: Left: Experimental results on N3DV dataset [li2022neural] showcase the effectiveness of our method, which reduces the storage requirement of 3DGStream [sun20243dgstream] by 159 $\times$, with enhanced visual quality. Right: Comparison with existing methods in storage and reconstruction fidelity. Hollow circles denote offline methods, while solid circles represent online methods.
@@ 1 Introduction
Reconstructing free-viewpoint video (FVV) from multi-view videos captured by cameras with known poses has attracted growing interest in the field of computer vision and graphics. FVV exhibits great potential as a next-generation visual medium that enables immersive and interactive experiences, with broad application in virtual and augmented reality (VR/AR) applications [chen2019overview].
@@
Recently, 3D Gaussian Splatting (3DGS) has become a promising method for FVV reconstruction, due to its significant advancements in real-time rendering and high-fidelity view synthesis. These approaches typically fall into two categories: 1) incorporating temporal function into Gaussian primitives and optimizing directly [yang2023real, luiten2024dynamic, li2024spacetime], and 2) applying a deformation field to capture the spatio-temporal transformations of canonical Gaussians [yang2024deformable, wu20244d, lu20243d, liu2025light4gs, bae2024per]. While these FVV reconstructions accurately represent dynamic scenes, they are trained in an offline manner and require transmitting the full set of reconstructed parameters prior to rendering.
@@
In contrast, by enabling per-frame training and progressive transmission, online FVV reconstruction allows immediate playback without the overhead of full-scene preloading. As a pioneer work, 3DGStream [sun20243dgstream] extends 3DGS to online FVV reconstruction using InstantNGP [muller2022instant] to model the geometric transformation frame-by-frame. While achieving impressive rendering speed, its structural constraint hinders the volumetric representation performance and degrades the visual quality. Building on this paradigm, subsequent works [girish2025queen, gaohicom] enhance model expressiveness through explicitly optimizing Gaussian attribute residuals, achieving competitive synthesis quality and higher robustness. However, the storage demands of these methods remain prohibitively high for real-time transmission, with reconstructed data typically exceeding 20 MB per second.
@@
In this paper, we aim to design a storage-efficient solution for FVV streaming that minimizes bandwidth requirements and enables real-time transmission. In online FVV reconstruction, since dynamic scenes contain a large proportion of static regions, the key to efficient reconstruction lies in motion modeling. Our first insight, therefore, is to only model the Gaussian attribute residuals in the motion regions, which eliminates the unnecessary updates in static regions. Building on motion modeling, we note that scene motion tends to be consistent, where Gaussian points associated with the same object typically exhibit the same or similar motion in dynamic scene representation. Our second insight, based on this observation, is to use a shared motion representation to model the attribute residuals with similar motion. This contrasts with existing online methods [sun20243dgstream, girish2025queen] that utilize point-wise strategy to update the attribute residuals in motion regions, and the result is motion redundancy elimination and more compact storage. Lastly, we exploit a key frame fine-tune strategy to handle the error accumulation brought by non-rigid motion and novel objects emergence.
@@
Specifically, to accomplish this, we propose a Compact Gaussian Streaming (ComGS) framework that leverages a set of keypoints (=200=200), significantly fewer than the full set of Gaussian points ($\approx 200K$), to holistically model motion regions at each timestep. ComGS begins with a motion-sensitive keypoints selection through a viewspace gradient difference strategy. This ensures that the selected keypoints are accurately positioned within motion regions and prevents redundant or incorrect modeling of static areas. Subsequently, we design an adaptive motion-driven mechanism that defines a keypoint-specific spatial influence field, with which neighboring Gaussian points can share the motion of the keypoint. Unlike conventional k-nearest neighbor (KNN) methods [huang2024sc, wan2024superpoint], the spatial influence field can accommodate the complexity and variability of motion structure in dynamic scenes, so that keypoints can more accurately drive the motion of the surrounding region. Finally, to mitigate error accumulation in a compact and effective manner, we propose an error-aware correction strategy for key frame reconstruction that selectively updates only those Gaussians with reconstruction errors.
@@
  • Experiments on two benchmark datasets show that the effectiveness of our method and its individual components. Our method achieves a compression ratio of $159\times$ over the 3DGStream and $14\times$ over state-of-the-art model QUEEN, enabling real-time transmission while preserving competitive reconstruction quality and rendering speed.
@@ 2.1 Dynamic Gaussian Splatting
Recently, 3D Gaussian Splatting (3DGS) [kerbl20233d, navaneet2023compact3d, fan2023lightgaussian, lee2024compact, wang2025end, papantonakis2024reducing] has attracted great attention in Free-viewpoint video (FVV) reconstruction for its high photorealistic performance and real-time rendering speed. Several works [yang2023real, luiten2024dynamic, li2024spacetime, xu2024representing, wang2025freetimegs] expand temporal variation as a function and optimize directly for modeling Gaussian attributes across frames. For instance, 4D Gaussian Splatting [yang2023real] incorporates time-conditioned 3D Gaussians and auxiliary components into 4D Gaussians, while ST-GS [li2024spacetime] models the transformation of structural attributes and opacity as a temporal function to represent scene motions. To support long FVVs representation, TGH [xu2024representing] introduces a multi-level hierarchy of 4D Gaussian primitives that exploits various degrees of temporal redundancy in dynamic scenes. While these time variant-based methods achieve superior rendering efficiency, they often suffer from prohibitive storage requirements. Other works [wu20244d, bae2024per, cho20244d, sun2025sdd, yan20244d] employ vanilla 3D Gaussians as a canonical space and a deformation field to represent the dynamic scene. In this category, 4D-GS [wu20244d] utilizes hexplanes [cao2023hexplane], six orthogonal planes, as latent embeddings and deliver them into a small MLP to deform temporal transformation of Gaussian points, achieving efficient computational complexity and lightweight storage requirement. Building upon this, GD-GS [bae2024per] further improves scene modeling accuracy by incorporating geometric priors, which provides a more structured and precise representation of dynamic scene. Among them, both SC-GS [huang2024sc] and SP-GS [wan2024superpoint] adopt sparse control points to control scene motion using a k-nearest neighbor (KNN) [peterson2009k] strategy for motion modeling. While these methods achieve notable improvements in computational efficiency and rendering speed, they are designed for offline FVV reconstruction and do not support frame-by-frame delivering. Additionally, motion-insensitive control point selection and scale-agnostic KNN motion modeling lead to redundant representation of static regions and reduced deformation accuracy in dynamic scenes. Our online method addresses these limitations by selecting keypoints from motion regions at each timestep and modeling motion with awareness of local motion scales, which enables more accurate and efficient modeling of online FVV.
@@ 2.2 Online Free-Viewpoint Video Reconstruction
Compared to the offline methods, online reconstruction enables FVV to be incrementally trained and transmitted in a per-frame manner, which allows users to preview or interact immediately with the video content. Leveraging the high-fidelity view synthesis capabilities of Neural Radiance Field (NeRF) [muller2022instant, mildenhall2021nerf, sun2022direct, chen2022tensorf, chen2023dictionary, fridovich2022plenoxels, barron2021mip, verbin2022ref], a set of studies have explored NeRF-based methods [li2022streaming, wang2024videorf, wang2023neural, wu2024tetrirf, song2023nerfplayer] for online FVV reconstruction, such as StreamRF [li2022streaming], VideoRF [wang2024videorf] and TeTriRF [wu2024tetrirf]. Despite advanced visual quality, NeRF-based methods are hindered by their limited rendering speeding of implicit structure, which limits their practical applications.
@@
With the utilization of 3DGS [kerbl20233d], 3DGStream [sun20243dgstream] introduces a hash-based MLP to encode the position and rotation transformation of Gaussian points at each frame, and designs an adaptive Gaussians addition strategy for novel objects across frames. Based on this paradigm, QUEEN [girish2025queen] proposes a Gaussian residual-based framework for model expressiveness enhancement and a learned quantization-sparsity framework for residuals compression. HiCoM [gaohicom] designs a hierarchical coherent motion mechanism to effectively capture and represent scene motion for fast and accurate training. To deploy into mobile device, V3 [wang2024v] presents a novel approach that compresses Gaussian attributes as a 2D video to facilitate hardware video codecs. IGS [yan2025instant] proposes a generalized anchor-driven Gaussian motion network that learns residuals with a singe step, achieving a significant improvement of training speed. Nevertheless, these methods face challenge in real-time transmission, due to their substantial storage requirements. This overhead mainly stems from redundant updates of static Gaussian points across frames, as well as repeated modeling of Gaussian points with similar motion. Our study exploits the locality and consistency of motion by leveraging motion-sensitive keypoints to adaptively drive motion regions, and this avoids redundant storage and transmission.
2.3 3D Gaussian Representation Compression
Despite 3DGS-based methods achieve impressive performance in novel view synthesis [kerbl20233d, lu2024scaffold], the massive size of Gaussian points hinder them for efficient storage and transmission. Several studies propose a variety of compression techniques for reducing the required storage, which can be categorized into either post-processing-based [fan2023lightgaussian, lee2024compact, wang2025end, niedermayr2024compressed, girish2025queen] or neural contextual coding-based methods [liu2025light4gs, chen2025hac, zhan2025cat, chen20254dgs, tang2025compressing]. Post-processing-based approaches include removing unimportant Gaussian points [fan2023lightgaussian, lee2024compact], pruning spherical harmonic coefficients [fan2023lightgaussian, wang2025end], and applying vector quantization [girish2025queen, niedermayr2024compressed] to compress the parameter representation. The latter methods utilize sophisticated entropy modeling to accurately predict probability distributions that exploit global context for compressing 3D Gaussian representation more effectively. In this paper, we focus on leveraging the locality and consistency of motion in dynamic scene and mitigating the redundancy reconstruction on static and similar motion regions, introducing a novel and more compact method for online FVV reconstruction.
@@
Figure 2: The overall pipeline of ComGS framework. (a) The reconstruction process starts from the first frame initialized using vanilla 3DGS [kerbl20233d]. Subsequent frames are organized into groups of frames (GoFs). For non-key frames, (b) we begins with a motion-sensitive keypoint selection using a viewspace gradient difference strategy, (c) and utilizes an adaptive motion-driven mechanism to control neighboring points motion. For key frames, (d) an error-aware correction strategy is introduced to mitigate the error accumulation across frames.
@@ 3.1 Preliminary
3DGS [kerbl20233d] models a 3D scene as a large amount of anisotropic 3D Gaussian points in world space as an explicit representation. The central position and geometric shape of each Gaussian point $i$ in world space are defined by a mean vector $\mu_{i}$ and covariance matrix $\Sigma_{i}$, mathematically represented as:
$G_{i}(x)=\text{exp}({-\frac{1}{2}(\mathbf{x}-\mu_{i})^{T}\Sigma_{i}^{-1}(\mathbf{x}-\mu_{i}))}$ (1)
For differentiable optimization, the covariance matrix $\Sigma_{i}$ is decoupled into a scaling matrix $S_{i}$ and a rotation matrix $R_{i}$. Each Gaussian point is characterized by its color $c_{i}$ and opacity $\sigma_{i}$. For novel view synthesis, the covariance matrix $\Sigma^{{}^{\prime}}_{i}$ in camera coordinate is given as:
$\Sigma^{{}^{\prime}}_{i}=\mathbf{J}\mathbf{W}\Sigma_{i}\mathbf{W}^{T}\mathbf{J}^{T}$ (2)
where $\mathbf{J}$ is the Jacobian of the affine approximation of the perspective projection and $\mathbf{W}$ represents the view transformation matrix mapping world coordinates.
During rendering, the Gaussian points are initially projected into viewing plane, and the final color $C$ can be obtained by $\alpha$-blending of the $N$ ordered 3D Gaussian points overlapping the pixel as:
$C=\sum_{i=1}^{N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$ (3)
where $\alpha_{i}$ represents the blending weight of the $i^{th}$ Gaussian point.
@@ 3.2 Motion-Sensitive Keypoint Selection
Thus, inspired by [girish2025queen], we propose a motion-sensitive keypoint selection based on viewspace gradient difference (Fig. 2 (b)). The core idea is to identify the dynamic Gaussian by the gradient change of rendering loss in inter-frames, and based on the gradient values, the $k$ Gaussian points with the largest gradient are selected as keypoints. Specifically, following the gradient computation in 3DGS [kerbl20233d], we compute gradients using the previous Gaussian positions $p_{t-1}$, the rendered images $\hat{I}_{t-1}$, the reconstruction loss $\mathcal{L}{{}_{recon}}$, and the ground-truth images $I_{t-1}$ and $I_{t}$:
$\mathcal{G}_{t-1}=\frac{\partial\mathcal{L}_{recon}^{t-1}}{\partial p_{t-1}},\ \mathcal{L}_{recon}^{t-1}=\mathcal{L}_{recon}(\hat{I}_{t-1},I_{t-1})$ (4)
$\mathcal{G}_{t}=\frac{\partial\mathcal{L}_{recon}^{t}}{\partial p_{t-1}},\ \mathcal{L}_{recon}^{t}=\mathcal{L}_{recon}(\hat{I}_{t-1},I_{t})$ (5)
Dynamic significance scores $\Delta\mathcal{G}_{t}\in\mathbb{R}^{N}$ (NN is the number of Gaussians) were calculated by means of absolute values of gradient differences: